# Global well-posedness for small data in a 3D temperature-velocity model with Dirichlet boundary noise

### v2

@@ 3.2. The remainder temperature equation
we deduce
$\Psi(\zeta_{1})-\Psi(\zeta_{2})=-(-\Delta)^{-\frac{s}{2}}\int_{0}^{t}(-\Delta)^{\gamma+\frac{s}{2}+\frac{1}{2}}e^{\left(t-r\right)\Delta}(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{r}\zeta_{r}\right)dr.$
Considering the norm, we obtain
(3.30) $\begin{split}\norm{\Psi(\zeta_{1})-\Psi(\zeta_{2})}_{X}&\leq\sup_{t\in[0,T_{1}]}\norm{\int_{0}^{t}(-\Delta)^{\gamma+\frac{s}{2}+\frac{1}{2}}e^{\left(t-r\right)\Delta}(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{r}\zeta_{r}\right)dr}_{L^{\frac{6}{5}}(\mathcal{D})}\\ &\leq\int_{0}^{T_{1}}\norm{(-\Delta)^{\gamma+\frac{s}{2}+\frac{1}{2}}e^{\left(T_{1}-r\right)\Delta}(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{r}\zeta_{r}\right)}_{L^{\frac{6}{5}}(\mathcal{D})}dr.\end{split}$
Using the same argument applied in Step 2 of the proof of Corollary 3.3, we get
(3.31) $\begin{split}\norm{\Psi(\zeta_{1})-\Psi(\zeta_{2})}_{X}&\leq\int_{0}^{T_{1}}\frac{c}{\left(T_{1}-r\right)^{\gamma+\frac{s}{2}+\frac{1}{2}}}\left\|(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{r}\zeta_{r}\right)\right\|_{L^{\frac{6}{5}}(\mathcal{D})}dr\\ &\leq\left(\int_{0}^{T_{1}}\left(\frac{c}{\left(T_{1}-r\right)^{\gamma+\frac{s}{2}+\frac{1}{2}}}\right)^{\frac{p}{p-\lambda-\delta_{u}}}dr\right)^{\frac{p-\lambda-\delta_{u}}{p}}\\ &\quad\cdot\left(\int_{0}^{T_{1}}\left\|(-\Delta)^{-\gamma-\frac{1}{2}}\,\mathrm{div}(u_{r}\zeta_{r})\right\|_{L^{\frac{6}{5}}(\mathcal{D})}^{\frac{p}{\lambda+\delta_{u}}}dr\right)^{\frac{\lambda+\delta_{u}}{p}}.\end{split}$
where $c=c(\gamma(s),s)=c(s)$. The first term is finite thanks to condition (3.26). The second term can be bounded as at the end of Step 3 in (3.29), to obtain
$\begin{split}\left(\int_{0}^{T_{1}}\left\|(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{r}\zeta_{r}\right)\right\|_{L^{\frac{6}{5}}(\mathcal{D})}^{\frac{p}{\lambda+\delta_{u}}}dr\right)^{\frac{\lambda+\delta_{u}}{p}}\leq C\sup_{t\in[0,T_{1}]}\norm{\zeta}_{L^{q}(\mathcal{D})}\norm{u}_{L^{\frac{p}{\lambda+\delta_{u}}}\left(0,T_{1};W^{\frac{1}{2}+\lambda,2}(\mathcal{D})\right)}.\end{split}$
Here, $C=C(\lambda(s),\gamma(s))=C(s)$ is independent of the time variable. In conclusion, we have shown
(3.32) $\norm{\Psi(\zeta_{1})-\Psi(\zeta_{2})}_{C\left(0,T_{1};W^{s,\frac{6}{5}}(\mathcal{D})\right)}\leq K\,\norm{u}_{L^{\frac{p}{\lambda+\delta_{u}}}\left(0,T_{1};W^{\frac{1}{2}+\lambda,2}(\mathcal{D})\right)}\,\norm{\zeta}_{C\left(0,T_{1};W^{s,\frac{6}{5}}(\mathcal{D})\right)},$
where we observe that $K$ is independent of the time variable. Thus, since $u\in L^{\frac{p}{\lambda+\delta_{u}}}\left(0,T;W^{\frac{1}{2}+\lambda,2}(\mathcal{D})\right)$, then $\Psi$ is a contraction from $X$ to itself if we choose $T_{1}>0$ sufficiently small. This gives the existence of a solution $\zeta\in C\left(0,T_{1};W^{s,\frac{6}{5}}(\mathcal{D})\right)$. If $T_{1}<T$, we can iteratively repeat the fixed point argument to get the solution of the thesis $\zeta\in C\left(0,T;W^{s,\frac{6}{5}}(\mathcal{D})\right)$ over the entire time interval $[0,T]$.
@@ 4.1. Stokes operator in weak setting, $\mathcal{H}^{\infty}$-calculus and maximal regularity
We start this section by recalling how the weak Stokes operator $A_{w}$ inherits the property of a bounded $\mathcal{H}^{\infty}$-calculus with $\mathcal{H}^{\infty}$-angle $\varphi^{\infty}_{A_{w}}=0$ from the (strong) Stokes operator $A$. This is obtained by following the approach in [25, Section 5].
@@
The first result that we recall is the maximal regularity of the weak Stokes operator, which is a consequence of the $\mathcal{H}^{\infty}$-calculus recalled above, see [24, Section II.4].
@@
Second, we recall the mixed derivative theorem, see e.g. [24, Corollary 4.5.10].
@@ 5. Well-posedness for the 3D-coupled temperature-velocity model with boundary noise
Fix $T>0$, $s\in[0,\frac{1}{2})$ and $p>4$. We proceed to verify that parameters satisfying the hypotheses exist. First, we choose $\gamma=\frac{1}{4}+\varepsilon_{\gamma}$ with $\varepsilon_{\gamma}>0$ small. Recall $a(s,\gamma)=\frac{3}{4}+\frac{s}{2}+\varepsilon_{\gamma}$. Since $s<1/2$, we have $a(s,\gamma)<1$ for small $\varepsilon_{\gamma}$. We require $\lambda_{\max}>\lambda_{\min}$, i.e., $p(1-a(s,\gamma))>\frac{1}{2}-s$. As $\varepsilon_{\gamma}\to 0$, this condition becomes
$p\left(1-\frac{3}{4}-\frac{s}{2}\right)>\frac{1}{2}-s\iff p\frac{1-2s}{4}>\frac{1-2s}{2}\iff p>2.$
Since we assumed $p>4>2$, the previous condition is satisfied. Next, we choose $\lambda\in(\lambda_{\min},\lambda_{\max})\cap(0,1)$ and define $\overline{\delta}_{\theta},\overline{\delta}_{u}$ as in the statement. We must show that the interval for $\delta_{u}$ is non-empty, i.e.,
$\max\{\delta_{\theta},\tfrac{1}{2}-s\}\leq\delta_{u}<\min\{\overline{\delta}_{u},1-\tfrac{2}{p}\}.$
We can always choose $\delta_{\theta}$ small enough such that $\delta_{\theta}\leq 1/2-s$. Thus, we need to ensure $\frac{1}{2}-s<\min\{\overline{\delta}_{u},1-\frac{2}{p}\}$. This is verified, up to choosing $\varepsilon_{\gamma}>0$ small enough, with the same arguments as in Remark 2.9. Finally, note that the condition $\delta_{u}<1-2/p$ implies $p>\frac{2}{1-\delta_{u}}$, which ensures the maximal regularity applicability in Theorem 2.5.
@@
Let $\tilde{\eta}=\tilde{\eta}(p,\delta_{u},T)>0$ and $M=M(p,\delta_{u},T)>0$ be the constants given by Theorem 2.5. Since $\delta_{u}\geq\max\{\delta_{\theta},\frac{1}{2}-s\}$, the embeddings
@@
$\tau^{\varepsilon}:=\inf\left\{t>0\ :\ \norm{Z^{\varepsilon}_{t}}_{H^{-2\alpha_{\theta}}(\mathcal{D})}>\frac{\eta}{8\widetilde{M}}\right\}$,
@@
To apply the Banach fixed point theorem, we first check that $\Gamma$ maps $\mathbb{B}_{1}$ into itself. We can use (2.8) by Corollary 2.3 since its hypothesis, namely $\delta_{\theta}\leq\overline{\delta}_{\theta}$ and $\delta_{u}<\overline{\delta}_{u}$ are satisfied; further, we have the embedding $\mathbb{E}_{\tau^{\varepsilon},p}^{\delta_{u}}\hookrightarrow L^{\frac{p}{\lambda+\delta_{u}}}(0,\tau^{\varepsilon};W^{\frac{1}{2}+\lambda,2}(\mathcal{D})),$ which holds for any $0<\lambda\leq 1-\delta_{u}$. Thus,
@@ Acknowledgments
G.D.S. acknowledges the support from the DFG project FOR 5528. M.L. is supported by the the Italian national inter-university PhD course in Sustainable Development and Climate change. M.L. produced this work while attending the PhD programme in PhD in Sustainable Development And Climate Change at the University School for Advanced Studies IUSS Pavia, Cycle XXXVIII, with the support of a scholarship financed by the Ministerial Decree no. 351 of 9th April 2022, based on the NRRP - funded by the European Union - NextGenerationEU - Mission 4 "Education and Research", Component 1 "Enhancement of the offer of educational services: from nurseries to universities” - Investment 3.4 “Advanced teaching and university skills”. The authors thank Franco Flandoli and Eliseo Luongo for the fruitful discussion on the topic.

### v3

@@ 3.2. The remainder temperature equation
we deduce that, for every $t\in[0,T_{1}]$,
$\Psi(\zeta_{1})(t)-\Psi(\zeta_{2})(t)=-(-\Delta)^{-\frac{s}{2}}\int_{0}^{t}(-\Delta)^{\gamma+\frac{s}{2}+\frac{1}{2}}e^{(t-r)\Delta}(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\bigl(u_{r}\zeta_{r}\bigr)\,dr.$
Set
$\sigma:=\gamma+\frac{s}{2}+\frac{1}{2},\qquad h(r):=(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}(u_{r}\zeta_{r}),$
and recall that $q:=\frac{p}{\lambda+\delta_{u}}$ and $q^{\prime}=\frac{q}{q-1}$. Using the smoothing estimate for the heat semigroup, we obtain for every $t\in[0,T_{1}]$
$\begin{split}\|\Psi(\zeta_{1})(t)-\Psi(\zeta_{2})(t)\|_{W^{s,\frac{6}{5}}(\mathcal{D})}&\leq C\int_{0}^{t}\|(-\Delta)^{\sigma}e^{(t-r)\Delta}h(r)\|_{L^{\frac{6}{5}}(\mathcal{D})}\,dr\\ &\leq C\int_{0}^{t}(t-r)^{-\sigma}\|h(r)\|_{L^{\frac{6}{5}}(\mathcal{D})}\,dr.\end{split}$
Applying Hölder’s inequality in time with exponents $q,q^{\prime}$, we get
$\|\Psi(\zeta_{1})(t)-\Psi(\zeta_{2})(t)\|_{W^{s,\frac{6}{5}}}$ $\leq C\left(\int_{0}^{t}(t-r)^{-\sigma q^{\prime}}\,dr\right)^{\frac{1}{q^{\prime}}}\left(\int_{0}^{t}\|h(r)\|_{L^{\frac{6}{5}}}^{q}\,dr\right)^{\frac{1}{q}}$
$\leq C\left(\int_{0}^{T_{1}}\tau^{-\sigma q^{\prime}}\,d\tau\right)^{\frac{1}{q^{\prime}}}\|h\|_{L^{q}(0,T_{1};L^{\frac{6}{5}}(\mathcal{D}))}.$
The integral in $\tau$ is finite thanks to condition (3.29), i.e. $\sigma q^{\prime}<1$. Hence,
(3.34) $\|\Psi(\zeta_{1})-\Psi(\zeta_{2})\|_{X}\leq C(\int_{0}^{T_{1}}\tau^{-\sigma q^{\prime}}\,d\tau)^{\frac{1}{q^{\prime}}}\|h\|_{L^{q}(0,T_{1};L^{\frac{6}{5}}(\mathcal{D}))}.$
Finally, repeating the argument of Step3 in (3.32)-(3.33) we obtain
$\|h\|_{L^{q}(0,T_{1};L^{\frac{6}{5}}(\mathcal{D}))}\leq C\|u\|_{L^{\frac{p}{\lambda+\delta_{u}}}\left(0,T_{1};W^{\frac{1}{2}+\lambda,2}(\mathcal{D})\right)}\|\zeta\|_{X}.$
Substituting into (3.34) yields
(3.35) $\|\Psi(\zeta_{1})-\Psi(\zeta_{2})\|_{X}\leq K(T_{1})\|u\|_{L^{\frac{p}{\lambda+\delta_{u}}}(0,T_{1};W^{\frac{1}{2}+\lambda,2}(\mathcal{D}))}\|\zeta_{1}-\zeta_{2}\|_{X},$
where
$K(T_{1}):=C\left(\int_{0}^{T_{1}}\tau^{-\sigma q^{\prime}}\,d\tau\right)^{\frac{1}{q^{\prime}}}.$
Since $u\in L^{\frac{p}{\lambda+\delta_{u}}}(0,T;W^{\frac{1}{2}+\lambda,2}(\mathcal{D}))$, we have $\|u\|_{L^{\frac{p}{\lambda+\delta_{u}}}(0,T_{1};W^{\frac{1}{2}+\lambda,2})}\to 0$ as $T_{1}\downarrow 0$, and moreover $K(T_{1})<\infty$ for every $T_{1}>0$ by $\sigma q^{\prime}<1$. Therefore, choosing $T_{1}>0$ sufficiently small makes the right-hand side of (3.35) strictly less than $\|\zeta_{1}-\zeta_{2}\|_{X}$, so $\Psi$ is a contraction on $X$. This yields a unique fixed point $\zeta\in C(0,T_{1};W^{s,\frac{6}{5}}(\mathcal{D}))$. If $T_{1}<T$, we may iterate the fixed point argument finitely many times to cover $[0,T]$. ∎
@@ 4.1. Stokes operator in weak setting, $\mathcal{H}^{\infty}$-calculus and maximal regularity
We start this section by recalling how the weak Stokes operator $A_{w}$ inherits the property of a bounded $\mathcal{H}^{\infty}$-calculus with $\mathcal{H}^{\infty}$-angle $\varphi^{\infty}_{A_{w}}=0$ from the (strong) Stokes operator $A$. This is obtained by following the approach in [27, Section 5].
@@
The first result that we recall is the maximal regularity of the weak Stokes operator, which is a consequence of the $\mathcal{H}^{\infty}$-calculus recalled above, see [26, Section II.4].
@@
Second, we recall the mixed derivative theorem, see e.g. [26, Corollary 4.5.10].
@@ 5. Well-posedness for the 3D-coupled temperature-velocity model with boundary noise
Fix $T>0$, $s\in(0,\tfrac{1}{2})$ and $p>4$. Choose, as in (2.8),
$\gamma:=\frac{1}{4}+\varepsilon_{\gamma},\qquad 0<\varepsilon_{\gamma}<\min\left\{\frac{1-2s}{4},\frac{(1-2s)(p-4)}{4p}\right\},$
and set
$a(s,\gamma):=\frac{1}{2}+\gamma+\frac{s}{2},\qquad\lambda_{\max}:=p(1-a(s,\gamma)),\qquad\lambda_{\min}:=\frac{1}{2}-s.$
Then $\lambda_{\max}>\lambda_{\min}$, thanks to the same computation as in (2.9). Hence the interval $(\lambda_{\min},\lambda_{\max})$ is non-empty. Thus, we may choose $\lambda\in(\lambda_{\min},\lambda_{\max})$. In particular, as in (2.11), we pick
$\lambda:=\frac{1}{2}-s+\varepsilon_{\lambda}\qquad 0<\varepsilon_{\lambda}<\min\{2s,\lambda_{\max}-(1-2s)\}.$
Lastly, as in (2.12), define the thresholds
$\overline{\delta}_{\theta}:=\min\left\{\lambda,2\gamma-\frac{1}{2}\right\},\qquad\overline{\delta}_{u}:=\min\left\{1-\lambda,\lambda_{\max}-\lambda\right\}.$
With this choice we have
$1-\lambda>\frac{1}{2}-s,\qquad\lambda_{\max}-\lambda>\frac{1}{2}-s,\qquad\text{hence}\qquad\overline{\delta}_{u}>\frac{1}{2}-s.$
In preparation for the application of Sobolev embeddings used below and Theorem 2.5, we now fix $\delta_{u}\in(0,\overline{\delta}_{u})$ and $0<\delta_{\theta}<\overline{\delta}_{\theta}$ satisfying the compatibility conditions (2.15), namely
$\delta_{u}\geq\max\left\{\delta_{\theta},\frac{1}{2}-s\right\},\qquad\delta_{u}<1-\frac{2}{p}.$
These constraints are consistent: indeed $p>4$ implies $1-\frac{2}{p}>\frac{1}{2}$ and, by the choice of $\lambda$, we have $\overline{\delta}_{u}>\frac{1}{2}-s$, hence one can pick $\delta_{u}$ in the non-empty interval $\bigl[\frac{1}{2}-s,\min\{\overline{\delta}_{u},1-\frac{2}{p}\}\bigr)$ and then choose any $\delta_{\theta}\in(0,\min\{\overline{\delta}_{\theta},\delta_{u}\})$. Finally, note that $\delta_{u}<1-\frac{2}{p}$ is equivalent to $p>\frac{2}{1-\delta_{u}}$, which is the integrability requirement needed to apply maximal regularity in Theorem 2.5.
@@
Let $\tilde{\eta}=\tilde{\eta}(p,\delta_{u},T)>0$ and $M=M(p,\delta_{u},T)>0$ be the constants given by Theorem 2.5. Since $\delta_{u}\geq\max\{\delta_{\theta},\frac{1}{2}-s\}$ thanks to the compatibility conditions (2.15), the embeddings
@@
$\tau^{\varepsilon}:=\inf\left\{t\in[0,T]\ :\ \norm{Z^{\varepsilon}_{t}}_{H^{-2\alpha_{\theta}}(\mathcal{D})}>\frac{\eta}{8\widetilde{M}}\right\}$,
@@
To apply the Banach fixed point theorem, we first check that $\Gamma$ maps $\mathbb{B}_{1}$ into itself. We can use (2.13) by Corollary 2.3 since its hypothesis, namely $\delta_{\theta}<\overline{\delta}_{\theta}$ and $\delta_{u}<\overline{\delta}_{u}$ are satisfied; further, we have the embedding $\mathbb{E}_{\tau^{\varepsilon},p}^{\delta_{u}}\hookrightarrow L^{\frac{p}{\lambda+\delta_{u}}}(0,\tau^{\varepsilon};W^{\frac{1}{2}+\lambda,2}(\mathcal{D})),$ which holds for our fixed $\lambda$ thanks to the assumption $\delta_{u}<1-\lambda$. Thus,
@@ Acknowledgments
We thank the two anonymous reviewers for their careful reading of the manuscript and for their valuable comments and suggestions, which helped improve the paper. G.D.S. acknowledges the support from the DFG project FOR 5528. M.L. is supported by the Italian national inter-university PhD course in Sustainable Development and Climate change. M.L. produced this work while attending the PhD programme in PhD in Sustainable Development And Climate Change at the University School for Advanced Studies IUSS Pavia, Cycle XXXVIII, with the support of a scholarship financed by the Ministerial Decree no. 351 of 9th April 2022, based on the NRRP - funded by the European Union - NextGenerationEU - Mission 4 "Education and Research", Component 1 "Enhancement of the offer of educational services: from nurseries to universities” - Investment 3.4 “Advanced teaching and university skills”. The authors thank Franco Flandoli and Eliseo Luongo for the fruitful discussion on the topic.